OpenStudy (anonymous):

how do you find the total positive area under the curve from x=0 to x=2 for y=e^x-2x?

OpenStudy (anonymous):

Just take the derivative from 0 to 2 \[\int\limits_{0}^{2} (e^x - 2x) dx\]

OpenStudy (anonymous):

i did, but this is a different question. It says i should solve for f(x) first. Its the third problem on the second page of this pdf

OpenStudy (anonymous):

In 1 a, you are solving for the integral of f(x), not f(x). The integral is what the answer to the last question you posted is, e^x - 2x So if I define g(x) = e^x - 2x then the definite integral of f(x) from a to b is g(b)-g(a) This will solve the next part (b).

OpenStudy (anonymous):

ok, i did part b, but dont understand what to do for part c